OpenStudy (hy123):

15^4x=4^x-4 Solve for x. Write the exact answer using the base-10 logarithms.

OpenStudy (hy123):

\[15^{4x}=4^{x-4}\]

ganeshie8 (ganeshie8):

\[\huge 15^{4x} = 4^{x-4}\] its like this ?

OpenStudy (hy123):

Yeah.

ganeshie8 (ganeshie8):

take log on both sides

OpenStudy (hy123):

I did that.

OpenStudy (hy123):

Wouldn't you just make it so they both change to 60 and then multiply the exponent by the number you used to get to 60?

ganeshie8 (ganeshie8):

\[\huge log\ m^n = nlog\ m\]

ganeshie8 (ganeshie8):

i see.. you trying to solve for x without using log ?

OpenStudy (hy123):

Well they want the exact answer using the base-10 log

ganeshie8 (ganeshie8):

question is asking to leave the answer in log terms.. so we can simply separate the x

OpenStudy (hy123):

So I'm going to guess they want it in log terms.

ganeshie8 (ganeshie8):

yeah... u r correct

OpenStudy (hy123):

Is it that simply? 4xlog15=x-4log4?

ganeshie8 (ganeshie8):

yup... u forgot the parantheses.. 4xlog15=(x-4)log4

OpenStudy (hy123):

So that's all they are asking for when they say solve for x?

OpenStudy (hy123):

Because they want me to slve for x while writing the answer in log terms.

ganeshie8 (ganeshie8):

yeah... but we still need to separate the x completely

OpenStudy (hy123):

Ahh okay lol.

OpenStudy (hy123):

Mind showing me the steps? I'm not really sure how to get there.

ganeshie8 (ganeshie8):

4xlog15=(x-4)log4

ganeshie8 (ganeshie8):

4xlog15 = xlog4 - 4log4

ganeshie8 (ganeshie8):

subtract xlog4 so that all x terms come one side

OpenStudy (hy123):

4xlog15-xlog4=-4log4

ganeshie8 (ganeshie8):

yep.. now factor x

OpenStudy (hy123):

x(4log15-log4)=-4log4?

ganeshie8 (ganeshie8):

yh..

OpenStudy (hy123):

So the answer is (-4log4)/(4log15-log4)?

ganeshie8 (ganeshie8):

yep.. thats it. if u dont like minus sign on top, reverse the order of subtraction on base..

OpenStudy (hy123):

Well i dont mind. Thanks very much.

ganeshie8 (ganeshie8):

welcome... good work :))